OpenStudy (anonymous):

Decided whether the statement are always, sometimes or never true. --> If 'a' and 'b' are rational numbers, then the product of ( a + √b) and its conjugate is a rational number.

OpenStudy (anonymous):

always since \[(a+\sqrt{b})(a-\sqrt{b})=a^2-b\]

OpenStudy (agreene):

oh, product of its conjugate.. didnt read that word... satellite wins... again >.<